log 10 2=a and log 10 3=b.if 3 x+2 =45.then value of 1-a/b

log102=a and log103=b.if 3x+2=45.then value of 1-a/b

Grade:12

2 Answers

Vikas TU
4 years ago
Dear student 
1-a/b = 1- log2/log3 
the value would be 0.37 …........................
Hope this will help 
Good Luck 
Devanarayanan V
4 years ago
log 2 = a
log 3 = b
1-a/b= 1-(log 2/log 3)
1- log32
Pls put the value and solve
Hope this was helpful